I need help with this word problem please
A forensic scientist uses the expression
72.6 + 2.5T to estimate the height in centimeters of a
female with a tibia of length T centimeters. If a female
skeleton has a tibia of length 32.4 cm, then what was the
height of the person? Find the length of your tibia in centimeters,
and use the expression from this exercise or the
previous exercise to estimate your height.

Answer by Mathtut(3670)

H(t)=height
:
H(t)=72.6+2.5t
:
H(32.4)=72.6+2.5(32.4)=153.6 cm
:
now they want you to take your height (in centimeters). remember 1 inch equals 2.54 centimeters. So lets take me for example I am 75 inches tall. so first I multiply that by 2.54 to find my height in centimeters. 75(2.54)=190.5 then I plug it into H(t) part of the equation and solve for t to find my tibia size
:
190.5=76.6+2.5t
:
113.9=2.5t
:
t=45.56 centimeters-length of my tibia. Now in inches that would be 45.56/2.54=17.93 inches.